OpenStudy (anonymous):

what is the solution of the following system? {-2x-y=1 {-4x-2y=1 A infinitely many solutions B no solution C (3,8) D (-3,-8)

OpenStudy (uri):

Did you try solving it?

OpenStudy (anonymous):

yes i did but i dont think i did it right

OpenStudy (uri):

Did you solve this using elimination method? Well,Show your work.

OpenStudy (anonymous):

okay well i did it by multiplying 2 from -2x-y=1 which would get -4x-2y=2 - { -4x-2y=-1 = -8x-4=3 -8x=7 x= -8/7

OpenStudy (uri):

I would say B cuz. -2x-y=1(multiply by -2 to make it positive) -4x-2y=1 So, 4x+2y=-1 -4x-2y=1 ======== Everything will be cancelled...so yeah.

OpenStudy (anonymous):

okay thank you very much

OpenStudy (uri):

Yw :)